determine if sequence is arithmetic, geometric, or neither. if arithmetic find common difference and the sum of the first n terms. if geometric find common ratio and sum of the first n terms
3333 32/4/8/16
Answer #1

Soli Given 3,3 3 4 3 8 3 TE ometric geom The given sequence is a Sequence Common ratio Cr) =ta t j = 3 2 2 x = 8x/ 3 3 y = 2So Sn: 3 (1 -66%) x 2 2 6 (14) rato Y = 1 / and So = 661-6?) Hence Common the terms Sum of